# Regression of Suspension Violin Modes in KAGRA O3GK Data with Kalman Filters

### v2

@@ Abstract
Suspension thermal modes in interferometric gravitational-wave detectors produce narrow, high-$Q$ spectral lines that can contaminate gravitational searches and bias parameter estimation. In KAGRA, cryogenic mirrors are held by thick suspension fibers, designed to sustain such a low-temperature environment, which may further affect inharmonicity modes, fiber dimensions, and mechanical behavior compared to typical interferometers. Remaining a prominent source of narrowband contamination, we implement a Kalman filter to model and track violin lines, building on the methodology introduced in [1], and apply subtraction to KAGRA O3GK data. Using gravitational-wave template injections, we validate that the subtraction preserves matched-filter SNR while effectively suppressing line power. Comparisons of power spectral densities and residual analyses confirm that the method removes deterministic line contributions without introducing waveform distortions. This approach provides a cleaner strain channel for searches and parameter estimation and will become increasingly important for future low-temperature detectors with higher-$Q$ suspensions, such as the Einstein Telescope.
@@ I Introduction
An adaptive filtering technique based on state-space modeling was pioneered in GW field by Finn et al. [1, 8]. By representing each violin mode as a damped harmonic oscillator in a discrete-time state-space framework, one can track its dynamics—amplitude and phase—in real time, predict its evolution, and subtract it from the strain channel while consistently accounting for both the physical process and measurement noise. In this work, we implement a Kalman filter to reduce suspension violin modes from KAGRA O3GK public strain data  [^1] https://gwosc.org/O3/O3GK/ (Accessed August 2025). The filter is applied to long stretches of data, including segments with injected gravitational-wave signals to validate its performance. We assess its impact on matched-filter SNR, power spectral densities (PSDs), and residual waveforms, and compare results across different scenarios, including validation tests using the LIGO 40 m interferometer. This study is the first demonstration of extracting suspension thermal noise at KAGRA in the cryogenic regime and shows that the technique effectively suppresses narrow spectral lines while preserving gravitational-wave signals.
@@ II Experimental Setup
Figure 1: Schematic of the KAGRA interferometer [3]. All mirrors with labels are suspended inside the vacuum tanks with four types of vibration isolation system. Different types of circles in the figure represent the different types of vibration isolation system. Vacuum tanks in front of the input and end test masses (depicted as dotted grey circles) contain narrow angle baffles and optical systems for photon calibrator. ITMX (Y): input test mass X (Y), ETMX (Y): end test mass X (Y), BS: beam splitter, PRM: power recycling mirror, SRM: signal recycling mirror, IMMT (OMMT): input (output) mode-matching telescope, IFI (OFI): input (output) Faraday isolator.
@@
Figure 2: Measured O3GK sensitivity at GPS 1270312960 [2]. Our focus will be on the study of the first three violin modes referring to the prominent suspension thermal noises.
@@ II.1 Suspension Design at KAGRA
KAGRA employs a multi-stage pendulum suspension system to isolate its test masses from seismic and thermal noise shown in Fig. 3. The final stage consists of a sapphire test mass suspended by sapphire fibers, operated at cryogenic temperatures to reduce thermal fluctuations. The high quality factor ($Q\gtrsim 10^{7}$) of these fibers results in narrow, long-lived resonances — known as violin modes — in the 170-186 Hz frequency range, along with higher harmonics.
Figure 3: Schematic of Type-A suspension [3]. The suspension is a multiple-stage pendulum with a total height of 13m. The Type-A tower is at room temperature while the cryopayload is cooled down to cryogenic temperature.
Fig. 4 shows the power spectral density (PSD) in the DARM channel. Each of the $N$ resonant peak in the PSD is modeled as a Lorentzian profile, overall fitted mode by mode, as the sum of Lorentzian distributions.
$S_{x}(f)=\sum_{i=1}^{N}\frac{G_{i}\,f_{0,i}^{2}\Gamma_{i}^{2}}{\left(f^{2}-f_{0,i}^{2}\right)^{2}+f^{2}\Gamma_{i}^{2}},$ (1)
@@
No background contribution is including in this fit. Also, for high-$Q$ resonances, the quality factor of each mode is given by
$Q_{i}=\frac{f_{0,i}}{\mathrm{\Gamma}_{i}}.$ (2)
@@
In cryogenic conditions, the coupling of thermal noise can significantly differ from that in room-temperature suspensions. This makes its suppression particularly relevant for KAGRA, especially given the discrepancy between theoretical expectations and experimental measurements reported in [2]. Furthermore, the violin-mode resonances do not align with exact integer multiples of the fundamental frequency. In particular, the second and third harmonics deviate from the usual harmonic scaling expected for a thin fiber, as exhibited by the LIGO40m setup in Fig. 15. This deviation illustrates that the sapphire suspension wires in KAGRA are subject to material rigidity effects [10], which give rise to inharmonicity, in close analogy to what has been observed in the LIGO and Virgo experiments. In the evaluation of three first-mode peaks, we observed that only 27, 30 and 31 peaks were identified shown in Fig. 4 and Table. 1. Faint or missing peaks hinder direct estimation of the $Q_{i}$ factor, while two closely spaced resonances may blend into a single feature, leading to an artificially inflated $Q$ value. This effect is examined in detail in the following section.
@@ II.2 Second order resonant effects
The CAD model of the mirror and its suspension wires is shown in Fig. 5. In addition to the fundamental violin modes of the suspension fibers along the beam axis, denoted $x$-axis, a coupling from higher-order resonances and other degrees of freedom - such as yaw, pitch, and roll rotations around the mirror’s center of mass - may introduce secondary spectral features. Further details can be found in Appendix C. In a first-order angular development, only the yaw contribution (Oz) is expected to be significant, since the lever arm is set by approximately half the mirror diameter along the y-axis for each suspension wire. The pitch degree of freedom is neglected here under the assumption that the attachment points of the fibers are coplanar in the $xOy$ plane; if the wires were not perfectly coplanar, vertical offsets would introduce additional pitch coupling. The roll degree of freedom is neglected at first order because displacements along the $x$-axis are largely insensitive to rotations around this axis.
@@
In practice, these second-order couplings can produce weak sidebands or nearby peaks in the power spectral density (PSD) at frequency offsets $\Delta f_{0}$, complicating the filtering process in addition to the longitudinal-like resonance strongly coupled with the arm length’s readout. Auxiliary peaks, sharing a common $Q_{i}$ value, typically have much lower amplitudes as they couple through geometric asymmetries. To illustrate such an effect, we computed a high-resolution PSD in Fig. 6 similar to that in Finn et al., highlighting the secondary lower-amplitude yaw-like peak candidates. Some lower amplitude peaks may be buried in the measurement noise or degenerated with other resonances, making them less visible in the PSD distribution. Since the measurement noise is lower for the second and third modes, it is easier to perform proper peak extraction in this frequency region; explaining why only 27 peaks are found in the first mode, as reported in Table 1.
Figure 6: High-resolution PSD computed using KAGRA data. Each of the longitudinal-direction modes is expected to have a companion mode at higher frequency due to yaw coupling—see the pairs highlighted in green, purple, and orange. The candidate shifts of about $0.45,\mathrm{Hz}$ are consistent with the theoretical lower bound estimated in Appendix C.
@@ III.1 Suspension thermal dynamics
The dynamics of a single violin mode are modeled as a damped harmonic oscillator with resonance frequency $\omega_{0}$ and quality factor $Q$, described by the equation of motion
$m\ddot{x}(t)+b\,\dot{x}(t)+k\,x(t)=0,$ (3)
@@
$\omega_{0}=\sqrt{\frac{k}{m}},\qquad Q=\frac{m\omega_{0}}{b}.$ (4)
@@
$\ddot{x}(t)+\frac{\omega_{0}}{Q}\,\dot{x}(t)+\omega_{0}^{2}x(t)=\dfrac{1}{m}\xi(t),$ (5)
@@
$\langle\xi(t)\rangle=0,\qquad\langle\xi(t)\,\xi(t^{\prime})\rangle=2\lambda k_{B}T\,\delta(t-t^{\prime}),$ (6)
@@ III.2 Slowly varying envelope approximation
$x(t)\approx\tilde{x}(t)\,e^{i\omega_{0}t}+\tilde{x}^{*}(t)\,e^{-i\omega_{0}t},$ (8)
where $\tilde{x}(t)$ evolves slowly compared to $e^{\pm i\omega_{0}t}$. ( $\ddot{x}\ll\dot{x}$ ) Substituting into the oscillator equation yields
$\frac{\omega_{0}}{Q}\,\dot{x}(t)+\omega_{0}^{2}x(t)=\xi(t),$ (9)
@@
$p_{\pm}=-\frac{\omega_{0}}{2Q}\pm i\omega_{0},$ (10)
@@ III.3 Data conditioning
$\psi(t)=x(t)\,e^{-i\omega_{c}t},\qquad\omega_{c}=2\pi f_{c},$ (11)
@@
$\bm{A}\approx\begin{pmatrix}-\dfrac{\omega_{0}}{2Q}&\,\omega_{c}-\omega_{0}\\ \omega_{0}-\omega_{c}&-\dfrac{\omega_{0}}{2Q}\end{pmatrix}$. (13)
@@ III.4 Observation and state-space representation
$\left\{\begin{aligned} \bm{\psi}_{k+1}&=\mathbf{A}^{\prime}\,\bm{\psi}_{k}+\mathbf{w}_{k},\\ \mathbf{z}_{k}&=\mathbf{C}\,\bm{\psi}_{k}+\mathbf{v}_{k},\end{aligned}\right$. (16)
$\mathbf{w}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{W}),\quad\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{V}),$ (17)
@@
  • $\mathbf{A^{\prime}}$ is the discrete-time state transition matrix, derived from the oscillator parameters $\omega_{c}$, $\omega_{0}$ and $Q$ for each mode,
@@
Some peaks lie within few mHz, it is necessary to extend the Kalman filter formalism along the lines of Finn et al. [1, 8]. A joint extraction is then possible by defining the new matrix elements as :
@@
$\mathbf{W}$ $=\mathrm{diag}(\mathbf{W}_{1},\dots,\mathbf{W}_{N}),$ $\mathbf{V}$ $\propto\mathbf{I}_{2}.$
@@
$\mathbf{W_{i}}=\sigma_{w,i}^{2}\mathbf{I}_{n},\quad\mathbf{V}=\sigma_{v}^{2}\mathbf{I}_{2}.$ (18)
@@
$\left\{\begin{aligned} \sigma_{w,i}^{2}&=\dfrac{\omega_{0,i}k_{B}T}{4Q^{2}\omega_{c}^{2}m\,\Delta f_{c}},\\[4.30554pt] \sigma_{v}^{2}&=\dfrac{1}{\Delta f_{c}}\,\dfrac{\bar{P}_{m}}{2},\end{aligned}\right$. (19)
These quantities are estimated from the physical suspension parameters $(Q,f_{0},k,b)$, which are strongly temperature dependent. In particular, cryogenic operation at KAGRA provides a unique opportunity to study how the suspension material properties evolve with temperature and, consequently, how they affect the noise budget. Measuring these dependencies would be highly valuable for improving violin-mode modeling and subtraction in future observing runs.
@@
$S_{F}(\omega\approx\omega_{0})\;=\;\bigl|H^{-1}(i\omega)\bigr|^{2}\,S_{x}(\omega),$ (20)
@@ IV.1 Sampling synthetic data
Figure 8: (a) Top: Amplitude (in dB) of the reconstructed state vector compared with the noisy true state vector including process noise. (b) Middle: Unwrapped phase (in rad) of the reconstructed state vector, with the linear trend $2\pi f_{0}t$ subtracted to align in time. (c) Bottom: Residual error distribution in logarithmic scale (relative symmetric error).
@@ IV.2 Reproduction of LIGO 40m noise suppression
As a validation step, we reproduce the violin-mode subtraction performance demonstrated in the LIGO 40 m interferometer [1]. The 40 m testbed provides a strong reference for validating our implementation, allowing direct comparison between our implementation and previous established results.
@@
This first reproduction confirms that our Kalman filter framework reproduces the suppression performance and residual characteristics reported in earlier work from 2001 [1]. A key feature is the strong reduction of stationary PSD noise in the vicinity of the resonant peaks after Kalman filtering. In our framework, higher-order resonances that lie very close to one another are already modeled and subtracted jointly, ensuring robust suppression. Looking ahead, we aim to refine the estimation of the process noise covariance $W$ and the measurement noise covariance $V$ by having a deeper understanding of the underlying hardware system, enabling more optimal suppression. As shown in Fig. 9, the filter also achieves strong suppression (on the order of 10 dB) of the 10th harmonic of the 60 Hz U.S. power grid. This occurs because both the grid harmonics and the violin resonances are governed by similar differential equations, leading the filter to treat them in a coupled manner. For optimal performance, such additional spectral lines should also be modeled and corrected jointly. However, since the focus of this paper is on the impact in cryogenic operation, their treatment is left to future work.

### v3

@@ Abstract
Suspension thermal modes in interferometric gravitational-wave detectors produce narrow, high-$Q$ spectral lines that can contaminate gravitational searches and bias parameter estimation. In KAGRA, cryogenic mirrors are held by thick suspension fibers, designed to sustain such a low-temperature environment, which may further affect inharmonicity modes, fiber dimensions, and mechanical behavior compared to typical interferometers. As these modes remain a prominent source of narrowband contamination, we implement a Kalman filter to model and track violin lines, building on the methodology introduced in [1], and apply subtraction to KAGRA O3GK data. Using gravitational-wave template injections, we validate that the subtraction preserves matched-filter SNR while effectively suppressing line power. Comparisons of power spectral densities and residual analyses confirm that the method removes deterministic line contributions without introducing waveform distortions. This approach provides a cleaner strain channel for searches and parameter estimation and will become increasingly important for future low-temperature detectors with higher-$Q$ suspensions, such as the Einstein Telescope.
@@ I Introduction
An adaptive filtering technique based on state-space modeling was pioneered in GW field by Finn et al. [1, 8]. By representing each violin mode as a damped harmonic oscillator in a discrete-time state-space framework, one can track its dynamics—amplitude and phase—in real time, predict its evolution, and subtract it from the strain channel while consistently accounting for both the physical process and measurement noise. In this work, we implement a Kalman filter to reduce suspension violin modes from KAGRA O3GK public strain data  [^1] https://gwosc.org/O3/O3GK/ (Accessed August 2025). The filter is applied to long stretches of data, including segments with injected gravitational-wave signals, to validate its performance. We assess its impact on matched-filter SNR, power spectral densities (PSDs), and residual waveforms, and compare results across different scenarios, including validation tests using the LIGO 40 m interferometer. This study is the first demonstration of extracting suspension thermal noise at KAGRA in the cryogenic regime and shows that the technique effectively suppresses narrow spectral lines while preserving gravitational-wave signals.
@@ II Experimental Setup
Figure 1: Schematic of the KAGRA interferometer [3]. All mirrors with labels are suspended inside the vacuum tanks with four types of vibration isolation systems. The different types of circles in the figure represent the various types of vibration isolation systems. Vacuum tanks in front of the input and end test masses (depicted as dotted grey circles) contain narrow-angle baffles and optical systems for the photon calibrator. ITMX (Y): input test mass X (Y), ETMX (Y): end test mass X (Y), BS: beam splitter, PRM: power recycling mirror, SRM: signal recycling mirror, IMMT (OMMT): input (output) mode-matching telescope, IFI (OFI): input (output) Faraday isolator.
@@
Figure 2: O3GK strain sensitivity at GPS 1270312960, highlighting the first three violin modes, which are the most prominent suspension thermal noise and the focus of this work.
@@ II.1 Suspension Design at KAGRA
KAGRA employs a multi-stage pendulum suspension system to isolate its test masses from seismic and thermal noise, shown in Fig. 3. The final stage consists of a sapphire test mass suspended by sapphire fibers, operated at cryogenic temperatures to reduce thermal fluctuations. The high quality factor ($Q\gtrsim 10^{7}$) of these fibers results in narrow, long-lived resonances — known as violin modes — in the 170-186 Hz frequency range, along with higher harmonics.
Figure 3: Schematic of Type-A suspension [3]. The suspension is a multiple-stage pendulum with a total height of 13m. The Type-A tower is at room temperature, while the cryopayload is cooled to cryogenic temperatures.
Fig. 4 shows the power spectral density (PSD) in the DARM channel. The PSD is modeled as a sum of separately fitted Lorentzian distributions, one for each of the $N$ identified peaks
$S_{x}(f)=\sum_{i=1}^{N}\frac{G_{i}\,f_{0,i}^{2}\Gamma_{i}^{2}}{\left(f^{2}-f_{0,i}^{2}\right)^{2}+f^{2}\Gamma_{i}^{2}}$ (1)
@@
No background contribution is included in this fit. Also, for high-$Q$ resonances, the quality factor of each mode is given by
$Q_{i}=\frac{f_{0,i}}{\mathrm{\Gamma}_{i}}$ (2)
@@
In cryogenic conditions, the coupling of thermal noise can significantly differ from that in room-temperature suspensions. This makes its suppression particularly relevant for KAGRA, especially given the discrepancy between theoretical expectations and experimental measurements reported in [2]. Furthermore, the violin-mode resonances do not align with exact integer multiples of the fundamental frequency. In particular, the second and third harmonics deviate from the usual harmonic scaling expected for a thin fiber, as exhibited by the LIGO40m setup in Fig. 15. This deviation illustrates that the sapphire suspension wires in KAGRA are subject to material rigidity effects [10], which give rise to inharmonicity, in close analogy to what has been observed in the LIGO and Virgo experiments. In the evaluation of the first three mode peaks, we observed that only 27, 30, and 31 peaks were identified, as shown in Fig. 4 and Table. 1. Faint or missing peaks hinder direct estimation of the $Q_{i}$ factor, while two closely spaced resonances may blend into a single feature, leading to an artificially inflated $Q$ value. This effect is examined in detail in the following section.
@@ II.2 Second order resonant effects
The CAD model of the mirror and its suspension wires is shown in Fig. 5. In addition to the fundamental violin modes of the suspension fibers along the beam axis, denoted $x$, a coupling from higher-order resonances and other degrees of freedom - such as yaw, pitch, and roll rotations around the mirror’s center of mass - may introduce secondary spectral features. Further details can be found in Appendix C. In a first-order angular development, only the yaw contribution (Oz) is expected to be significant, since the lever arm is set by approximately half the mirror diameter along the y-axis for each suspension wire. The pitch degree of freedom is neglected here under the assumption that the attachment points of the fibers are coplanar in the $xOy$ plane; if the wires were not perfectly coplanar, vertical offsets would introduce additional pitch coupling. The roll degree of freedom is neglected at first order because displacements along the $x$-axis are largely insensitive to rotations around this axis.
@@
In practice, these second-order couplings can produce weak sidebands or nearby peaks in the power spectral density (PSD) at frequency offsets $\Delta f_{0}$, complicating the filtering process in addition to the longitudinal-like resonance strongly coupled with the arm length’s readout. Auxiliary peaks, sharing a common $Q_{i}$ value, typically have much lower amplitudes as they couple through geometric asymmetries. To illustrate such an effect, we computed a high-resolution PSD in Fig. 6 similar to that in Finn et al., highlighting the secondary lower-amplitude yaw-like peak candidates. Some lower amplitude peaks may be buried in the measurement noise or degenerate with other resonances, making them less visible in the PSD distribution. Since the measurement noise is lower for the second and third modes, it is easier to perform proper peak extraction in this frequency region; explaining why only 27 peaks are found in the first mode, as reported in Table 1.
Figure 6: High-resolution PSD computed using KAGRA data. Each of the longitudinal-direction modes is expected to have a companion mode at a higher frequency due to yaw coupling—see the pairs highlighted in green, purple, and orange. The candidate shifts of about $0.45\,\mathrm{Hz}$ are consistent with the theoretical lower bound estimated in Appendix C.
@@ III.1 Suspension thermal dynamics
The dynamics of a single violin mode are modeled as a damped harmonic oscillator with resonance frequency $\omega_{0}$ and quality factor $Q$, described by the equation of motion.
$m\ddot{x}(t)+b\,\dot{x}(t)+k\,x(t)=0$ (3)
@@
$\omega_{0}=\sqrt{\frac{k}{m}},\qquad Q=\frac{m\omega_{0}}{b}$ (4)
@@
$\ddot{x}(t)+\frac{\omega_{0}}{Q}\,\dot{x}(t)+\omega_{0}^{2}x(t)=\dfrac{1}{m}\xi(t)$ (5)
@@
$\langle\xi(t)\rangle=0,\qquad\langle\xi(t)\,\xi(t^{\prime})\rangle=2\lambda k_{B}T\,\delta(t-t^{\prime})$ (6)
@@ III.2 Slowly varying envelope approximation
$x(t)\approx\tilde{x}(t)\,e^{i\omega_{0}t}+\tilde{x}^{*}(t)\,e^{-i\omega_{0}t}$ (8)
where $\tilde{x}(t)$ evolves slowly compared to $e^{\pm i\omega_{0}t}$ ( $\ddot{x}\ll\dot{x}$ ). Substituting into the oscillator equation yields
$\frac{\omega_{0}}{Q}\,\dot{x}(t)+\omega_{0}^{2}x(t)=\xi(t)$ (9)
@@
$p_{\pm}=-\frac{\omega_{0}}{2Q}\pm i\omega_{0}$ (10)
@@ III.3 Data conditioning
$\psi(t)=x(t)\,e^{-i\omega_{c}t},\qquad\omega_{c}=2\pi f_{c}$ (11)
@@
$\bm{A}\approx\begin{pmatrix}-\dfrac{\omega_{0}}{2Q}&\,\omega_{c}-\omega_{0}\\ \omega_{0}-\omega_{c}&-\dfrac{\omega_{0}}{2Q}\end{pmatrix}$ (13)
@@ III.4 Observation and state-space representation
$\left\{\begin{aligned} \bm{\psi}_{k+1}&=\mathbf{A}^{\prime}\,\bm{\psi}_{k}+\mathbf{w}_{k}\\ \mathbf{z}_{k}&=\mathbf{C}\,\bm{\psi}_{k}+\mathbf{v}_{k}\end{aligned}\right$. (16)
$\mathbf{w}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{W}),\quad\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{V})$ (17)
@@
  • $\mathbf{A^{\prime}}$ is the discrete-time state transition matrix, derived from the oscillator parameters $\omega_{c}$, $\omega_{0}$, and $Q$ for each mode,
@@
Some peaks lie within a few mHz; it is therefore necessary to apply a joint Kalman filter extraction along the lines of Finn et al. [1, 8]. Such extraction is performed by defining the new matrix elements as :
@@
$\mathbf{W}$ $=\mathrm{diag}(\mathbf{W}_{1},\dots,\mathbf{W}_{N}),$ $\mathbf{V}$ $\propto\mathbf{I}_{2}$
@@
$\mathbf{W_{i}}=\sigma_{w,i}^{2}\mathbf{I}_{n},\quad\mathbf{V}=\sigma_{v}^{2}\mathbf{I}_{2}$ (18)
@@
$\left\{\begin{aligned} \sigma_{w,i}^{2}&=\dfrac{\omega_{0,i}k_{B}T}{4Q^{2}\omega_{c}^{2}m\,\Delta f_{c}}\\[4.30554pt] \sigma_{v}^{2}&=\dfrac{1}{\Delta f_{c}}\,\dfrac{\bar{P}_{m}}{2}\end{aligned}\right$. (19)
These quantities are estimated from the physical suspension parameters $(Q,f_{0},k,b)$, which are strongly temperature dependent. In particular, cryogenic operation at KAGRA presents a unique opportunity to investigate how the properties of the suspension material evolve with temperature and, consequently, how they impact the noise budget. Measuring these dependencies would be highly valuable for improving violin-mode modeling and subtraction in future observing runs.
@@
$S_{F}(\omega\approx\omega_{0})\;=\;\bigl|H^{-1}(i\omega)\bigr|^{2}\,S_{x}(\omega)$ (20)
@@ IV.1 Sampling synthetic data
Figure 8: (a) Top: Amplitude (in dB) of the reconstructed state vector compared with the noisy true state vector, including process noise. (b) Middle: Unwrapped phase (in rad) of the reconstructed state vector, with the linear trend $2\pi f_{0}t$ subtracted to align in time. (c) Bottom: Residual error distribution in logarithmic scale (relative symmetric error).
@@ IV.2 Reproduction of LIGO 40m noise suppression
As a validation step, we reproduce the violin-mode subtraction performance demonstrated in the LIGO 40 m interferometer [1]. The 40 m testbed provides a strong reference for validating our implementation, allowing direct comparison between our implementation and previously established results.
@@
This first reproduction confirms that our Kalman filter framework reproduces the suppression performance and residual characteristics reported in earlier work from 2001 [1]. A key feature is the strong reduction of stationary PSD noise in the vicinity of the resonant peaks after Kalman filtering. In our framework, higher-order resonances that lie very close to one another are already modeled and subtracted jointly, ensuring robust suppression. Looking ahead, we aim to refine the estimation of the process noise covariance $W$ and the measurement noise covariance $V$ by having a deeper understanding of the underlying hardware system, enabling more optimal suppression. As shown in Fig. 9, the filter also achieves strong suppression (on the order of 10 dB) of the 10th harmonic of the 60 Hz U.S. power grid. This occurs because both the grid harmonics and the violin resonances are governed by similar differential equations, leading the filter to treat them in a coupled manner. For optimal performance, such additional spectral lines should also be modeled and corrected jointly. However, since the focus of this paper is on the impact of cryogenic operation, their treatment is left to future work.